the perimeter of a rectangle is 65 inches. The ratio of the lenth to the width is 4:1. what is the lenth and width
The perimeter is 24, the ratio is 2:1, what is the lenth and width

Answer by mananth(16946)

The perimeter = 65 inches
1/2 perimeter = 32.5 inches
x/y = 4:1
x=4y
x+y =32.5
Plug the value of x in the equation.
4y+y= 32.5
5y= 32.5
y=6.5
So x = 4*6.5 = 26
The dimensions ar 6.5 by 26
Check
26+26+6.5+6.5= 65